OpenStudy (anonymous):

hey is 10 a perfect square?

OpenStudy (raffle_snaffle):

No

OpenStudy (anonymous):

No

OpenStudy (anonymous):

There is no integer square root of 10.

OpenStudy (zzr0ck3r):

bellas last question bella_rose "hey why is everyone on here an idiot"

OpenStudy (zzr0ck3r):

maybe we should lie to you....I say \[\sqrt{10}\times \sqrt{10}=10\] and thus yes 10 is a perfect square

OpenStudy (anonymous):

haha ^^